For what real values of a is x^2 + ax + 25 the square of a binomial? If you find more than one, then list your values in increasing order, separated by commas.

Answer by Boreal(15235)

a=10
x^2+10x+25=(x+5)^2
a=-10
x^2-10x+25=(x-5)^2
(-10, 10)